TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-04-27 14:46:00
subject: 4\14 Pt 1 ISS On Orbit Status 14-04-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

15 Apr 2003

ISS On-Orbit Status 14 Apr 2003

Part 1 of 2

All ISS systems continue to function nominally, except as noted
previously or below.  Week 20 is underway.  Also:  Today is Day 142
for Expedition 6 (since STS-113 launch on 11/23/02).

The crew started preparations for their return on Soyuz TMA-1/5S on
5/4, focusing today on handover procedures for the Russian segment
(RS).  [The task involved review of the crew handover (RPS)
procedures book and the incorporation of new data that the Expedition
7 crew will need.] 

As part of the preparations, the crew completed the scheduled fit
check of the "Kazbeks", the three contoured shock-absorbing seats in
the Soyuz descent capsule (SA).  This required them to don their
Sokol pressure suits, take their seats and measure the gap between
the top of the head and the top edge of the structure facing the head
with a ruler.  The results were reported to MCC-M.   [The Kazbek-Us
are designed to withstand g-loads during launch and orbital insertion
as well as during reentry and brake-rocket-assisted landing.  Each
seat has two positions: cocked (armed) and noncocked. In the cocked
position, they are raised to allow the shock absorbers to function
during touchdown.  The fit check assures that the crew, whose bodies
gain in length during longer-term stay in zero-G, will still be
adequately protected by the seat liners for their touchdown in
Kazakhstan on 5/4.] 

In the Service Module (SM), FE-1 Nikolai Budarin replaced the TA765B
ROM (read-only memory) storage units in three TA251M local
temperature sensor switch devices (commutators, LKT1A2, LKT1B2,
LKT1V2) with new spares.  There are 11 of these LKTs behind various
SM panels which receive, convert and transmit temperature sensor data
via the local analog telemetry switch (LKA) to the database.  [Before
exchanging the TA765B boxes, Budarin pulled their BITS2-12 onboard
telemetry system connections, after the ground had powered the BITS
off.  RS systems have been powered off during the BITS component
replacement.  All BITS connectors were later to be remated for
activation and checkout of the installation by the ground.] 

FE-2/SO Don Pettit hade another highly successful day with the MSG
(Microgravity Science Glovebox), today running test 24 of the InSPACE
(Investigating the Structure of Paramagnetic Aggregates from
Colloidal Emulsions) experiment.

Target areas for the Russian Uragan earth imaging program on
Nikolai's task list for today, cloud cover permitting, were the
Angara River and Irkutsk, Ulan-Ude and the Selenga River, Kolka
glacier in  the north of the Kazbek mountains, the Chirkeisk
reservoir, the hydrological situation on the Rivers Ishim (at Astana,
near the Kazakhstan landing site for Soyuz 5S), Irtysh (near
Pavlodar), Ob (near Biisk), Maly and Verkhny Yenisei (near Kyzyl). 

CDR Ken Bowersox worked in the Joint Airlock, returning its
configuration to pre-EVA conditions.  [This involved buttoning up the
EMU suits, checking switch settings, and stowing EVA hardware as
required.] 

In a tagup with Don Pettit, Huntsville/POC (Payload Operations
Center) specialists discussed the newly developed procedures for the
repair of the ARCTIC freezer/refrigerator.  The crew's comments will
be incorporated in the final version of the work-around procedures by
end of this week.  As requested by the crew, the work will be placed
on the "job jar" task list, and POC will work with the timeliners to
get the repair scheduled before the end of this increment.  [The
ARCTIC-1 has been inoperative due to two failed thermo-electric
cooling units since last December.  The procedure would restore some
capability to the system by bypassing the units.  ARCTIC's cooling
capability would be reduced, but tests are needed to show by how
much.  If functionality can be reasonably restored, new research
activities are being looked at which might be flown up on Progress.]

The crew performed the six-month maintenance of the TVIS (treadmill
with vibration isolation and stabilization).  [The three-hour task
required removal of the TVIS from the "pit" in the SM floor.  The
work, which is part of the yearly scheduled TVIS maintenance
activities, was followed by a one-hour physical exercise session on
the treadmill, serving as a checkout.] 

Budarin completed the regular daily maintenance of the SOZh life
support system in the SM, while Bowersox prepared the daily inventory
update file for downlink to the IMS (inventory management system)
database and performed the regular status checkup of autonomous
payloads in the Lab (PCG-STES010, CGBA).

The VTR-2 (video tape recorder #2) playback problem reported before
(4/11/03) turned out to be just a video cable issue, which has been
cleared meanwhile.

The EarthKAM system at the Lab window was deactivated and stowed.
The next session is scheduled to start on 4/25 (Friday), to be going
on during the Soyuz 6S docked phase.  EK images can be viewed on the
Internet at
http://datasystem.earthkam.ucsd.edu/cgi-bin/datasys/ek_images_station

In Part 1 of the IFM (in-flight maintenance) on the cardiac
defibrillator, Bowersox determined that the defibrillator is working
OK.  Part 2 showed that the device's PDIM (power data interface
module) is currently not charging its batteries.  Troubleshooting
continues, and the defibrillator is cleared for nominal operations,
as long as batteries are available (which may have to be replaced if
recharge cannot be restored). 

Budarin completed his regular daily 5-min. inspection of the BIO-5
Rasteniya-2/Lada-2 ("Plants-2") plant growth experiment.

All crewmembers performed their regular daily physical exercise
program on RED, TVIS and VELO with load trainer.

At 11:50am EDT, amateur radio operator Ken Bowersox chatted with
students of Lounsberry Hollow Middle School in the NW corner of New
Jersey in a rural community.  [The school has 947 fifth and sixth
graders, and the school district is the largest in the county.]

The predicted conjunction with the TRMM satellite (Tropical Rainfall
Measuring Mission, object #25063), reported on 4/11, has passed
without requiring avoidance action, after the satellite was
maneuvered by the Goddard Space Flight Center (GSFC).

Three old on-orbit HDDs (hard disk drives) have been reconfigured for
PCS (portable computer system) use and are now available as spares.
A fourth HDD can also be utilized if required.

The laptop of EXPRESS 1 (ER1) has developed a problem and is powered
off. ER1 itself continues to work nominally, and there is no issue
with lack of memory.

At today's General Designers Review (GDR) at RSC-Energia in Moscow,
Energia experts confirmed the Soyuz 6S' readiness for launch on
4/26.  The Soyuz vehicle will be sent to the fueling station on 4/16
(Wednesday), starting irreversible operations.  The Expedition 7 crew
will depart from Star City to Baikonur on 4/20.

Today's CEO (crew earth observations) targets, including
near-vertical targets from the city target list due to the current
LVLH attitude, were Lima, Peru (nadir pass; ESC [electronic still
camera]), Alexandria, Egypt (nadir pass; ESC), Tucson, Arizona (nadir
pass; ESC), and Albuquerque, New Mexico (nadir pass; ESC). CEO images
can be viewed at the websites http://eol.jsc.nasa.gov and
http://earthobservatory.nasa.gov

 - Continued -

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.